# Determination of the stably free cancellation property for orders

### v2

@@ 7.10. Choosing $\mathfrak{g}$ first
Recall that in §7.1 and §7.4, respectively, we first chose $\mathfrak{f}$ to be a proper full two-sided ideal of $\mathcal{M}$ that is contained in $\Lambda$ and then set $\mathfrak{g}:=\mathfrak{f}\cap\mathcal{O}_{C}$. In some applications, it is convenient to work the other way round: first choose a proper full ideal $\mathfrak{g}$ of $\mathcal{O}_{C}$ such that $\mathfrak{g}\mathcal{M}\subseteq\Lambda$ and then set $\mathfrak{f}:=\mathfrak{g}\mathcal{M}$. The following lemma shows that the latter approach is consistent with the former.
@@
Let $\mathfrak{g}$ be a proper full ideal of $\mathcal{O}_{C}$ such that $\mathfrak{f}:=\mathfrak{g}\mathcal{M}\subseteq\Lambda$. Then $\mathfrak{f}$ is a proper full two-sided ideal of $\mathcal{M}$ and $\mathfrak{f}\cap\mathcal{O}_{C}=\mathfrak{g}$.
@@
The first claim is clear. To prove the second claim, we can and do work componentwise and hence assume without loss of generality that $A=\mathrm{Mat}_{n}(D)$ for some $n\geq 1$ and some division algebra $D$ containing $K$. Let $F=Z(D)$ and note that $F=C$.
Let $\Delta$ be any maximal $\mathcal{O}_{F}$-order in $D$. Then $\Delta\cap F=\mathcal{O}_{F}$. Thus if $\alpha\delta\in\mathcal{O}_{F}$ for some $\delta\in\Delta$ and some $\alpha\in\mathcal{O}_{F}$ with $\alpha\neq 0$ then $\delta=\alpha^{-1}(\alpha\delta)\in\Delta\cap F=\mathcal{O}_{F}$. This implies that as an $\mathcal{O}_{F}$-module, $\Delta/\mathcal{O}_{F}$ is torsion-free and hence projective. Thus the canonical short exact sequence of $\mathcal{O}_{F}$-modules
$0\longrightarrow\mathcal{O}_{F}\longrightarrow\Delta\longrightarrow\Delta/\mathcal{O}_{F}\longrightarrow 0$
splits, which gives an isomorphism of $\mathcal{O}_{F}$-lattices $\Delta\cong\mathcal{O}_{F}\oplus\Delta/\mathcal{O}_{F}$. Abusing notation slightly, it follows that
$\mathfrak{g}\Delta\cap\mathcal{O}_{F}=\left(\mathfrak{g}\mathcal{O}_{F}\oplus\mathfrak{g}(\Delta/\mathcal{O}_{F})\right)\cap\mathcal{O}_{F}=\mathfrak{g}.$ (7.12)
If $n=1$ then $\mathcal{M}$ is a maximal $\mathcal{O}_{F}$-order in $D=A$ and so taking $\Delta=\mathcal{M}$ in equation (7.12) gives the desired result.
Suppose that $n\geq 2$ and again let $\Delta$ be any maximal $\mathcal{O}_{F}$-order in $D$. For each right ideal $\mathfrak{a}$ of $\Delta$, let $\Delta^{\prime}=\mathcal{O}_{l}(\mathfrak{a}):=\{x\in D\mid x\mathfrak{a}\subseteq\mathfrak{a}\}$, and let
$\mathcal{M}_{\mathfrak{a},n}=\left(\begin{array}[]{cccc}\Delta&\ldots&\Delta&\mathfrak{a}^{-1}\\ \vdots&\ddots&\vdots&\vdots\\ \Delta&\ldots&\Delta&\mathfrak{a}^{-1}\\ \mathfrak{a}&\ldots&\mathfrak{a}&\Delta^{\prime}\end{array}\right)$
denote the ring of all $n\times n$ matrices $(x_{ij})$ where $x_{11}$ ranges over all elements of $\Delta$, …, $x_{1n}$ ranges over all elements of $\mathfrak{a}^{-1}$, and so on. By [Rei03, (27.6)], $\mathcal{M}_{\mathfrak{a},n}$ is a maximal $\mathcal{O}_{F}$-order in $A$ and there exists a choice of $\mathfrak{a}$ such that there is an isomorphism of $\mathcal{O}_{F}$-orders $\theta:\mathcal{M}_{\mathfrak{a},n}\cong\mathcal{M}$. Since $\Delta$ was arbitrary in (7.12), we have $\mathfrak{g}\Delta\cap\mathcal{O}_{F}=\mathfrak{g}=\mathfrak{g}\Delta^{\prime}\cap\mathcal{O}_{F}$. Thus it follows easily from the explicit description of $\mathcal{M}_{\mathfrak{a},n}$ that we have $\mathfrak{g}\mathcal{M}_{\mathfrak{a},n}\cap\mathcal{O}_{F}=\mathfrak{g}$. Since $\theta(\mathcal{O}_{F})=\mathcal{O}_{F}$, $\theta(\mathfrak{g})=\mathfrak{g}$ and $\theta(\mathfrak{g}\mathcal{M}_{\mathfrak{a},n})=\mathfrak{g}\mathcal{M}$, we conclude that $\mathfrak{g}\mathcal{M}\cap\mathcal{O}_{F}=\mathfrak{g}$, as desired. ∎
@@ 8.1. Test lattices
The following lemma determines the localisations of certain test lattices.
@@
Suppose that $\mathfrak{g}$ satisfies Assumption (G). Let $\beta\in\mathcal{M}\cap A^{\times}$ such that $\mathcal{M}\beta+\mathfrak{f}=\mathcal{M}$. Let $\mathfrak{p}$ be a maximal ideal of $\mathcal{O}$. Then
@@
Now suppose that $\Lambda_{\mathfrak{p}}\neq\mathcal{M}_{\mathfrak{p}}$. Then $\mathfrak{f}_{\mathfrak{p}}\neq\mathcal{M}_{\mathfrak{p}}$ and so by Assumption (G) and Lemma 7.9 we have that $\mathfrak{f}_{\mathfrak{P}}\subseteq\mathrm{rad}(\mathcal{M}_{\mathfrak{P}})$ for all maximal ideals $\mathfrak{P}$ of $\mathcal{O}_{C}$ with $\mathfrak{P}\mid\mathfrak{p}$. Thus
@@
Suppose that $\mathfrak{g}$ satisfies Assumption (G). Let $\overline{\beta}\in\left(\mathcal{M}/\mathfrak{f}\right)^{\times}$. Then for every lift $\beta\in\mathcal{M}\cap A^{\times}$ we have that
@@
Assertion (v) is just the definition of $\tilde{\nu}$ (7.9). Assertion (vi) follows from (ii), (iii) and (iv) together with Theorem 7.10. We have that $X_{\beta}$ is stably free over $\Lambda$ if and only if $[X_{\beta}]_{\mathrm{st}}=0$ and by (7.10) we have $\ker(\mu)=\tilde{\nu}((\Lambda/\mathfrak{f})^{\times})$. Therefore assertion (vii) now follows from (vi). Finally, assertion (viii) follows from (iii) and (iv) together with Corollary 7.3. ∎
@@
Suppose that $\mathfrak{g}$ satisfies Assumption (G). Let $X$ be a locally free left ideal of $\Lambda$. Suppose that $X+\mathfrak{f}=\Lambda$ and that there exists $\beta\in\mathcal{M}\cap A^{\times}$ such that $\mathcal{M}X=\mathcal{M}\beta$. Then $\overline{\beta}\in(\mathcal{M}/\mathfrak{f})^{\times}$ and $[X_{\beta}]_{\mathrm{st}}=[X]_{\mathrm{st}}$ in $\mathrm{Cl}(\Lambda)$.
@@
We have $\mathcal{M}=\mathcal{M}(X+\mathfrak{f})=\mathcal{M}X+\mathfrak{f}=\mathcal{M}\beta+\mathfrak{f}$, and so $\overline{\beta}\in(\mathcal{M}/\mathfrak{f})^{\times}$. Then by Theorem 7.10 and Proposition 8.3 (vi) we have $[X]_{\mathrm{st}}=\mu(\tilde{\nu}(\overline{\beta}))=[X_{\beta}]_{\mathrm{st}}$ in $\mathrm{Cl}(\Lambda)$. ∎
@@
Suppose that $\mathfrak{g}$ satisfies Assumption (G). Let $\beta_{1},\beta_{2}\in\mathcal{M}\cap A^{\times}$ such that $\overline{\beta_{1}},\overline{\beta_{2}}\in(\mathcal{M}/\mathfrak{f})^{\times}$. Suppose that $\overline{\beta_{1}},\overline{\beta_{2}}$ have the same image in ×\$\mathcal{M}^{\times}\backslash(\mathcal{M}/\mathfrak{f})^{\times}{}{/}{}(\Lambda/\mathfrak{f})^{\times}$. Then the following statements hold:
@@
Suppose that $\mathcal{M}$ has SFC and that $\mathfrak{g}$ satisfies Assumption (G). Let $S\subseteq\mathcal{M}\cap A^{\times}$ be a set of representatives of $(\mathcal{M}/\mathfrak{f})^{\times}$. Let $T=\{\beta\in S\colon X_{\beta}\text{ is stably free over }\Lambda\}$. Then $\Lambda$ has SFC if and only if $X_{\beta}$ is free over $\Lambda$ for every $\beta\in T$.
@@
Recall from §3.1 that in order to show that $\Lambda$ has SFC, it suffices to show that every stably free $\Lambda$-lattice of rank $1$ is in fact free. So suppose that $X$ is a stably free $\Lambda$-lattice of rank $1$, that is, $X\in g(\Lambda)$ and $[X]_{\mathrm{st}}=0$ in $\mathrm{Cl}(\Lambda)$. By Lemma 8.6 we can and do suppose without loss of generality that $X$ is a locally free left ideal of $\Lambda$ such that $X+\mathfrak{f}=\Lambda$. Moreover, $\mathcal{M}X\cong\mathcal{M}\otimes_{\Lambda}X$ is stably free of rank $1$ over $\mathcal{M}$. Since $\mathcal{M}$ has SFC by hypothesis, there exists $\beta_{1}\in\mathcal{M}\cap A^{\times}$ such that $\mathcal{M}X=\mathcal{M}\beta_{1}$. Then by Corollary 8.4 we have $\overline{\beta_{1}}\in(\mathcal{M}/\mathfrak{f})^{\times}$ and $[X_{\beta_{1}}]_{\mathrm{st}}=[X]_{\mathrm{st}}=0$. Hence by definition of $S$, there exists $\beta_{2}\in S$ such that $\overline{\beta_{2}}=\overline{\beta_{1}}$. Then $[X_{\beta_{2}}]_{\mathrm{st}}=[X_{\beta_{1}}]_{\mathrm{st}}=0$ by Proposition 8.3 (vi) and so in fact $\beta_{2}\in T$. Thus $X_{\beta_{2}}$ is free over $\Lambda$ by the assumption on $T$. By Proposition 8.3 (viii) this implies that there exists $u\in\mathcal{M}^{\times}$ such that $\pi(\overline{\beta_{2}})=\pi(\overline{u})$. Since $\pi(\overline{\beta_{1}})=\pi(\overline{\beta_{2}})$, we have $\pi(\overline{\beta_{1}})=\pi(\overline{u})$. Thus $X$ is free over $\Lambda$ by Corollary 7.3. ∎
@@
Suppose that $\mathcal{M}$ has SFC and that $\mathfrak{g}$ satisfies Assumption (G). Let $S\subseteq\mathcal{M}\cap A^{\times}$ be a set of representatives of ×\$\mathcal{M}^{\times}\backslash(\mathcal{M}/\mathfrak{f})^{\times}{}{/}{}(\Lambda/\mathfrak{f})^{\times}$. Let
@@ 8.2. An algorithm for determining SFC
The following algorithm determines whether an order has SFC or not. However, it is only practical for orders in algebras of ‘small’ dimension and in many situations, it is often much faster to use Algorithm 9.1 or 10.3.
@@
    Compute the centre $C$ of $A$ and the decomposition of $A$ into simple $K$-algebras.
@@
    Adjust $\mathfrak{f}$ (and thus $\mathfrak{g}$) to ensure that $\mathfrak{g}$ satisfies Assumption (G).
@@
In step (3), the correctness of an output $\mathsf{false}$ follows from Corollary 3.2 and the correctness of an output $\mathsf{true}$ is clear. If step (4) is reached then $\Lambda$ must be non-maximal. Note that $\mathfrak{f}$ as computed in step (4) is a two-sided ideal of $\mathcal{M}$ such that $\mathfrak{f}\mathcal{M}\subseteq\Lambda$, and so $\mathfrak{f}$ must be proper. By Proposition 8.3 (v),(vii) we have $T=\{\beta\in S\colon X_{\beta}\text{ is stably free over }\Lambda\}$. Thus in step (11), the correctness of the output follows from Theorem 8.7. ∎
@@
Step (6): First note that for a given ideal $\mathfrak{g}$ there are only finitely many maximal ideals $\mathfrak{p}$ of $\mathcal{O}$ such that $\mathfrak{g}_{\mathfrak{p}}\neq\mathcal{O}_{C,\mathfrak{p}}$. If $\mathfrak{g}=\mathfrak{f}\cap\mathcal{O}_{C}$ fails Assumption (G) for $\mathfrak{P}\mid\mathfrak{p}$, then we replace $\mathfrak{f}$ by $\mathfrak{P}\mathfrak{f}$ and observe that $(\mathfrak{P}\mathfrak{f}\cap\mathcal{O}_{C})_{\mathfrak{P}}=\mathfrak{P}\mathfrak{f}_{\mathfrak{P}}\cap\mathcal{O}_{C,\mathfrak{P}}\subseteq\mathfrak{P}\left(\mathfrak{f}_{\mathfrak{P}}\cap\mathcal{O}_{C,\mathfrak{P}}\right)\subseteq\mathfrak{P}\mathcal{O}_{C,\mathfrak{P}}$.
@@
Step (8): In the notation of §7.7, compute $\infty_{i}$ for $i=1,\ldots,r$ using [NS09, Theorem 3.5]. The structure of the finitely generated abelian group $I_{\mathfrak{g}}/P_{\mathfrak{g}}^{+}$ can then be computed component-wise using [Coh00, Algorithm 4.3.1]. Using the output of Step (7) and the definition of $\tilde{\nu}$, then determine generators of $\tilde{\nu}((\Lambda/\mathfrak{f})^{\times})$.
@@
Step (11): Given $\beta\in T$, [BJ11, Algorithm 3.1] (see also [HJ20, Algorithm 4.1]) can be used to determine whether $X_{\beta}$ is free over $\Lambda$ (though stated for orders in group algebras, this algorithm can also be applied to orders in arbitrary finite-dimensional semisimple algebras as explained in [BJ11, Remark 4.9]). Note that this algorithm requires certain hypotheses, which ensure that one can find the preimage of the image of the canonical map ×$\mathcal{M}^{\times}\longrightarrow(\mathcal{M}/\mathfrak{f})^{\times}$, as well as a generator of $\mathcal{M}X_{\beta}$ as an $\mathcal{M}$-module. Since generators of $\mathcal{M}^{\times}$ can in principle be found using the algorithms of [BCNS15], and since $\mathcal{M}X_{\beta}=\mathcal{M}\beta$ by Proposition 8.3 (iii), these hypotheses are in fact unnecessary in this situation. Note that we also give an alternative freeness testing algorithm in §12 that may be more efficient under certain circumstances. ∎
@@
    If necessary, adjust $\mathfrak{g}$ to ensure that it satisfies Assumption (G).
@@
    Compute a set of representatives $S\subseteq\mathcal{M}\cap A^{\times}$ of ×\$\mathcal{M}^{\times}\backslash(\mathcal{M}/\mathfrak{f})^{\times}{}{/}{}(\Lambda/\mathfrak{f})^{\times}$.

### v3

@@ 7.10. Choosing $\mathfrak{g}$ first
Recall that in §7.1 and §7.4, respectively, we first chose $\mathfrak{f}$ to be a full two-sided ideal of $\mathcal{M}$ that is contained in $\Lambda$ and then set $\mathfrak{g}:=\mathfrak{f}\cap\mathcal{O}_{C}$. In some applications, it is convenient to work the other way round: first choose a full ideal $\mathfrak{g}$ of $\mathcal{O}_{C}$ such that $\mathfrak{g}\mathcal{M}\subseteq\Lambda$ and then set $\mathfrak{f}:=\mathfrak{g}\mathcal{M}$. The following lemma shows that the latter approach is consistent with the former.
@@
Let $\mathfrak{g}$ be a full ideal of $\mathcal{O}_{C}$ such that $\mathfrak{f}:=\mathfrak{g}\mathcal{M}\subseteq\Lambda$. Then $\mathfrak{f}$ is a full two-sided ideal of $\mathcal{M}$ and $\mathfrak{f}\cap\mathcal{O}_{C}=\mathfrak{g}$.
@@
The first claim is clear. For the second claim, it suffices to show that $\mathfrak{f}_{i}\cap\mathcal{O}_{K_{i}}=\mathfrak{g}_{i}$ for each $i=1,\ldots,r$. So we fix $i\in\{1,\ldots,r\}$, let $L:=K_{i}$ and replace $\mathfrak{g}_{i}$ by $\mathfrak{g}$, $\mathfrak{f}_{i}$ by $\mathfrak{f}$, $\mathcal{M}_{i}$ by $\mathcal{M}$ and $\mathcal{O}_{C}$ by $\mathcal{O}_{L}$. By [Rei03, Theorem 4.21], it suffices to show that $(\mathfrak{f}\cap\mathcal{O}_{L})_{\mathfrak{P}}=\mathfrak{g}_{\mathfrak{P}}$ for each maximal ideal $\mathfrak{P}$ of $\mathcal{O}_{L}$. Fix such an ideal $\mathfrak{P}$. Then $\mathcal{O}_{L,\mathfrak{P}}$ is a discrete valuation ring, and so we have $\mathfrak{g}_{\mathfrak{P}}=g_{\mathfrak{P}}\mathcal{O}_{L,\mathfrak{P}}$ for some $g_{\mathfrak{P}}\in\mathcal{O}_{L,\mathfrak{P}}$. Then we have
$(\mathfrak{f}\cap\mathcal{O}_{L})_{\mathfrak{P}}$ $=\mathfrak{f}_{\mathfrak{P}}\cap\mathcal{O}_{L,\mathfrak{P}}=\mathfrak{g}_{\mathfrak{P}}\mathcal{M}_{\mathfrak{P}}\cap\mathcal{O}_{L,\mathfrak{P}}=g_{\mathfrak{P}}\mathcal{M}_{\mathfrak{P}}\cap\mathcal{O}_{L,\mathfrak{P}}$
$=g_{\mathfrak{P}}\mathcal{M}_{\mathfrak{P}}\cap Z(\mathcal{M}_{\mathfrak{P}})=g_{\mathfrak{P}}Z(\mathcal{M}_{\mathfrak{P}})=g_{\mathfrak{P}}\mathcal{O}_{L,\mathfrak{P}}=\mathfrak{g}_{\mathfrak{P}}.$
The authors are grateful to an anonymous referee for suggesting this proof. ∎
@@ 8.1. Test lattices
The following lemma determines the localizations of certain test lattices.
@@
Suppose that $\mathfrak{g}$ is locally radical. Let $\beta\in\mathcal{M}\cap A^{\times}$ such that $\mathcal{M}\beta+\mathfrak{f}=\mathcal{M}$. Let $\mathfrak{p}$ be a maximal ideal of $\mathcal{O}$. Then
@@
Now suppose that $\Lambda_{\mathfrak{p}}\neq\mathcal{M}_{\mathfrak{p}}$. Then $\mathfrak{f}_{\mathfrak{p}}\neq\mathcal{M}_{\mathfrak{p}}$ and so by the assumption that $\mathfrak{g}$ is locally radical and Lemma 7.10 we have that $\mathfrak{f}_{\mathfrak{P}}\subseteq\mathrm{rad}(\mathcal{M}_{\mathfrak{P}})$ for all maximal ideals $\mathfrak{P}$ of $\mathcal{O}_{C}$ with $\mathfrak{P}\mid\mathfrak{p}$. Thus
@@
Suppose that $\mathfrak{g}$ is locally radical. Let $\overline{\beta}\in\left(\mathcal{M}/\mathfrak{f}\right)^{\times}$. Then for every lift $\beta\in\mathcal{M}\cap A^{\times}$ we have that
@@
Assertion (v) is just the definition of $\tilde{\nu}$ (7.6). Assertion (vi) follows from (ii), (iii) and (iv) together with Theorem 7.11. We have that $X_{\beta}$ is stably free over $\Lambda$ if and only if $X_{\beta}]_{\mathrm{st}}=0$ and by (7.7) we have $\ker(\mu)=\tilde{\nu}((\Lambda/\mathfrak{f})^{\times})$. Therefore assertion (vii) now follows from (vi). Finally, assertion (viii) follows from (iii) and (iv) together with Corollary 7.3. ∎
@@
Suppose that $\mathfrak{g}$ is locally radical. Let $X$ be a locally free left ideal of $\Lambda$. Suppose that $X+\mathfrak{f}=\Lambda$ and that there exists $\beta\in\mathcal{M}\cap A^{\times}$ such that $\mathcal{M}X=\mathcal{M}\beta$. Then $\overline{\beta}\in(\mathcal{M}/\mathfrak{f})^{\times}$ and $X_{\beta}]_{\mathrm{st}}=[X]_{\mathrm{st}}$ in $\mathrm{Cl}(\Lambda)$.
@@
We have $\mathcal{M}=\mathcal{M}(X+\mathfrak{f})=\mathcal{M}X+\mathfrak{f}=\mathcal{M}\beta+\mathfrak{f}$, and so $\overline{\beta}\in(\mathcal{M}/\mathfrak{f})^{\times}$. Then by Theorem 7.11 and Proposition 8.3 (vi) we have $X]_{\mathrm{st}}=\mu(\tilde{\nu}(\overline{\beta}))=[X_{\beta}]_{\mathrm{st}}$ in $\mathrm{Cl}(\Lambda)$. ∎
@@
Suppose that $\mathfrak{g}$ is locally radical. Let $\beta_{1},\beta_{2}\in\mathcal{M}\cap A^{\times}$ such that $\overline{\beta_{1}},\overline{\beta_{2}}\in(\mathcal{M}/\mathfrak{f})^{\times}$. Suppose that $\overline{\beta_{1}},\overline{\beta_{2}}$ have the same image in $\mathcal{M}^{\times}\backslash(\mathcal{M}/\mathfrak{f})^{\times}{}{/}{}(\Lambda/\mathfrak{f})^{\times}$. Then the following statements hold:
@@
Suppose that $\mathcal{M}$ has SFC and that $\mathfrak{g}$ is locally radical. Let $S\subseteq\mathcal{M}\cap A^{\times}$ be a set of representatives of $(\mathcal{M}/\mathfrak{f})^{\times}$. Let $T=\{\beta\in S\colon X_{\beta}\text{ is stably free over }\Lambda\}$. Then $\Lambda$ has SFC if and only if $X_{\beta}$ is free over $\Lambda$ for every $\beta\in T$.
@@
Recall from §3.1 that in order to show that $\Lambda$ has SFC, it suffices to show that every stably free $\Lambda$-lattice of rank $1$ is in fact free. So suppose that $X$ is a stably free $\Lambda$-lattice of rank $1$, that is, $X\in g(\Lambda)$ and $X]_{\mathrm{st}}=0$ in $\mathrm{Cl}(\Lambda)$. By Lemma 8.6 we can and do suppose without loss of generality that $X$ is a locally free left ideal of $\Lambda$ such that $X+\mathfrak{f}=\Lambda$. Moreover, $\mathcal{M}X\cong\mathcal{M}\otimes_{\Lambda}X$ is stably free of rank $1$ over $\mathcal{M}$. Since $\mathcal{M}$ has SFC by hypothesis, there exists $\beta_{1}\in\mathcal{M}\cap A^{\times}$ such that $\mathcal{M}X=\mathcal{M}\beta_{1}$. Then by Corollary 8.4 we have $\overline{\beta_{1}}\in(\mathcal{M}/\mathfrak{f})^{\times}$ and $X_{\beta_{1}}]_{\mathrm{st}}=[X]_{\mathrm{st}}=0$. Hence by definition of $S$, there exists $\beta_{2}\in S$ such that $\overline{\beta_{2}}=\overline{\beta_{1}}$. Then $X_{\beta_{2}}]_{\mathrm{st}}=[X_{\beta_{1}}]_{\mathrm{st}}=0$ by Proposition 8.3 (vi) and so in fact $\beta_{2}\in T$. Thus $X_{\beta_{2}}$ is free over $\Lambda$ by the assumption on $T$. By Proposition 8.3 (viii) this implies that there exists $u\in\mathcal{M}^{\times}$ such that $\pi(\overline{\beta_{2}})=\pi(\overline{u})$. Since $\pi(\overline{\beta_{1}})=\pi(\overline{\beta_{2}})$, we have $\pi(\overline{\beta_{1}})=\pi(\overline{u})$. Thus $X$ is free over $\Lambda$ by Corollary 7.3. ∎
@@
Suppose that $\mathcal{M}$ has SFC and that $\mathfrak{g}$ is locally radical. Let $S\subseteq\mathcal{M}\cap A^{\times}$ be a set of representatives of $\mathcal{M}^{\times}\backslash(\mathcal{M}/\mathfrak{f})^{\times}{}{/}{}(\Lambda/\mathfrak{f})^{\times}$. Let
@@ 8.2. An algorithm for determining SFC
The following algorithm determines whether an order has SFC. However, it is only practical for orders in algebras of ‘small’ dimension and in many situations, it is often much faster to use Algorithm 9.1 or 10.3.
@@
    Compute the center $C$ of $A$ and the decomposition of $A$ into simple $K$-algebras.
@@
    Adjust $\mathfrak{f}$ (and thus $\mathfrak{g}$) to ensure that $\mathfrak{g}$ is locally radical.
@@
In step (3), the correctness of an output $\mathsf{false}$ follows from Corollary 3.2 and the correctness of an output $\mathsf{true}$ is clear. By Proposition 8.3 (v),(vii) we have $T=\{\beta\in S\colon X_{\beta}\text{ is stably free over }\Lambda\}$. Thus in step (11), the correctness of the output follows from Theorem 8.7. ∎
@@
Step (6): First note that for a given ideal $\mathfrak{g}$ there are only finitely many maximal ideals $\mathfrak{p}$ of $\mathcal{O}$ such that $\mathfrak{g}_{\mathfrak{p}}\neq\mathcal{O}_{C,\mathfrak{p}}$. If $\mathfrak{g}=\mathfrak{f}\cap\mathcal{O}_{C}$ is not locally radical then there exists such a $\mathfrak{p}$ and a maximal ideal $\mathfrak{P}$ of $\mathcal{O}_{C}$ with $\mathfrak{P}\mid\mathfrak{p}$ such that $\mathfrak{g}_{\mathfrak{P}}\not\subseteq\mathfrak{P}\mathcal{O}_{C,\mathfrak{P}}$. For each such $\mathfrak{P}$, we replace $\mathfrak{f}$ by $\mathfrak{P}\mathfrak{f}$ and observe that $(\mathfrak{P}\mathfrak{f}\cap\mathcal{O}_{C})_{\mathfrak{P}}=\mathfrak{P}\mathfrak{f}_{\mathfrak{P}}\cap\mathcal{O}_{C,\mathfrak{P}}\subseteq\mathfrak{P}\left(\mathfrak{f}_{\mathfrak{P}}\cap\mathcal{O}_{C,\mathfrak{P}}\right)\subseteq\mathfrak{P}\mathcal{O}_{C,\mathfrak{P}}$.
@@
Step (8): In the notation of §7.7, compute $\infty_{i}$ for $i=1,\ldots,r$ using [NS09, Theorem 3.5]. The structure of the finitely generated abelian group $I_{\mathfrak{g}}/P_{\mathfrak{g}}^{+}$ can then be computed componentwise using [Coh00, Algorithm 4.3.1]. Using the output of Step (7) and the definition of $\tilde{\nu}$, then determine generators of $\tilde{\nu}((\Lambda/\mathfrak{f})^{\times})$.
@@
Step (11): Given $\beta\in T$, [BJ11, Algorithm 3.1] (see also [HJ20, Algorithm 4.1]) can be used to determine whether $X_{\beta}$ is free over $\Lambda$ (though stated for orders in group algebras, this algorithm can also be applied to orders in arbitrary finite-dimensional semisimple algebras as explained in [BJ11, Remark 4.9]). Note that this algorithm requires certain hypotheses, which ensure that one can find the preimage of the image of the canonical map $\mathcal{M}^{\times}\longrightarrow(\mathcal{M}/\mathfrak{f})^{\times}$, as well as a generator of $\mathcal{M}X_{\beta}$ as an $\mathcal{M}$-module. Since generators of $\mathcal{M}^{\times}$ can in principle be found using the algorithms of [BCNS15], and since $\mathcal{M}X_{\beta}=\mathcal{M}\beta$ by Proposition 8.3 (iii), these hypotheses are in fact unnecessary in this situation. Note that we also give an alternative freeness testing algorithm in §12 that may be more efficient under certain circumstances. ∎
@@
    If necessary, adjust $\mathfrak{g}$ to ensure that it is locally radical.
@@
    Compute a set of representatives $S\subseteq\mathcal{M}\cap A^{\times}$ of $\mathcal{M}^{\times}\backslash(\mathcal{M}/\mathfrak{f})^{\times}{}{/}{}(\Lambda/\mathfrak{f})^{\times}$.